OpenStudy (naveenbhatia1312):

Find the exact solution to the equation. (5 points) 7-log_2(x+5) = 6 x = 3 x = 7 x = -3 x = -6

OpenStudy (sweetburger):

does that say log base 2?

OpenStudy (naveenbhatia1312):

yes

OpenStudy (naveenbhatia1312):

so the answer is 7?

OpenStudy (naveenbhatia1312):

@sweetburger

OpenStudy (mathmale):

7-log_2(x+5) = 6 As before, subtract 7 from both sides, OR, alternatively, Subtract 6 from both sides, obtaining \[1-\log_{2} (x-5)=0\]

OpenStudy (mathmale):

Move the log term to the right side:\[1=\log_{2} (x-5)\]... How would you now solve for x-5, and then for x alone? Note that you must use 2 as base here, and the corresponding exponential function must also have base 2.

OpenStudy (naveenbhatia1312):

I have no idea...

OpenStudy (mathmale):

Are you working on a test, a quiz or homework? What have you done with logarithmic functions in the near past? exponential functions?

OpenStudy (naveenbhatia1312):

homework? And I am fairly new to this

OpenStudy (mathmale):

Can you say how y=log x and y=10^x are related?

OpenStudy (naveenbhatia1312):

log is based by the number 10

OpenStudy (mathmale):

Those two functions have the same base, yes. One is the inverse of the other. What does "inverse function" mean to you?

OpenStudy (naveenbhatia1312):

inverse means opposite

OpenStudy (mathmale):

In this context "inverse" has a particular meaning. if the log function is used as the INPUT to the base 10 expo. function, we must obtain the following: \[10^{\log x}=x\]

OpenStudy (mathmale):

Also, because these two are inverse functions, \[\log 10^x=x\]

OpenStudy (mathmale):

If you look carefully and think about this, you'll see that one of these functions "undoes" the other. If I start with x and take the log to the base 10 of that, and then use the result as the exponent of 10, the result will be x, the quantity from which I started.

OpenStudy (naveenbhatia1312):

how does this relate to the problem

OpenStudy (mathmale):

Given\[1=\log_{2} (x-5)\]

OpenStudy (mathmale):

We want to find (1) x-5, and (2) x. What is the base of this log system, and how do you know that?

OpenStudy (mathmale):

This is a direct application of what we were discussing just before now.

OpenStudy (naveenbhatia1312):

base is x I believe

OpenStudy (mathmale):

No, the base is 2. Earlier I used 10 as the base. The same property applies here:\[2^{\log_{2}x }=x\] because these inverse functions "undo" each other.

OpenStudy (mathmale):

Given \[1=\log_{2} (x-5)\]

OpenStudy (mathmale):

how would you solve for x-5?

OpenStudy (mathmale):

Recall: here the base is 2.

OpenStudy (naveenbhatia1312):

?? can you do it cause I am not understanding

OpenStudy (mathmale):

Write '2' on both sides of a new equation. Then use '1' as the exponent of the '2' on the left. Can you show me your work?

OpenStudy (naveenbhatia1312):

I cant not able to

OpenStudy (mathmale):

Naveen, I have already led you through some examples. Have y ou not studied log functions and exponential functions before?

OpenStudy (naveenbhatia1312):

Im just confused and tired. 5 am where I live

OpenStudy (mathmale):



OpenStudy (mathmale):